OpenStudy (anonymous):

1. Write a proof for the biconditional using the inverse and contrapositive. x ≥ 4 if and only if 3x ≥ 12. 2. Write a proof for the statement. If a quadrilateral is a square, then the quadrilateral is a rectangle Please need help

OpenStudy (anonymous):

\[x\geq 4\implies 3\times x\geq 3\times 4\] is one direction

OpenStudy (anonymous):

\[3x\geq 12\implies \frac{3x}{3}\geq \frac{12}{3}\] goes the other way